# Deep Transformer Network for Monocular Pose Estimation of Ship-Based UAV

### v1

@@ 4.4 Model Inference with Bayesian Fusion
where $I_{0},I_{1}$ correspond to the modified Bessel function of the first kind, and $s\geq 0$ is a scalar that can be obtained numerically from $d$ [21]. The value of the above probability for varying $d$ and $\theta$ is illustrated at Figure 11. For example, when $d=0.999$, the estimation error is less than $5.07^{\circ}$ with the probability of $0.95$.
Figure 11: Degree of confidence in the attitude estimate
@@ 5.1 Model Training with Synthetic Data
The TNN-MO model was implemented using the PyTorch framework. The model was trained for 350 epochs with the batch size of 48, hyperparameter $\gamma=10$, leveraging the AdamW optimizer [22], using a 20GB Multi-Instance GPU (MIG) partition from NVIDIA A100-PCIE-40GB GPU. The learning rate was initially set to $10^{-4}$ for the first 200 epochs, after which it was reduced to $10^{-5}$. Gradient clipping was also employed, limiting the maximum gradient norm to 0.1. The resulting training curve is presented at Figure 12.
Figure 12: Training loss over epochs
@@ 5.2 Model Testing with Synthetic Data
Figure 13: TNN-MO model tested over synthetic data: for object classes with the confidence $o_{j}>0.9$, the key points and the base frame are illustrated
Figure 14: TNN-SO model (top) vs TNN-MO model (bottom) tested over long range synthetic data: for TNN-MO model, object classes with the confidence $o_{j}>0.9$, the key points are illustrated in colors. Compared to the TNN-SO model, the TNN-MO model exhibits the grater accuracy greater accuracy over a longer range.
@@ 6.1 Data Collection System (DCS)
The hardware configuration to capture a validation dataset is as follows. A data collection system (DCS) that was developed for ship air wake measurements [10] is augmented with a camera. It is attached to an octocoper unmanned aerial vehicle that is manually controlled to fly around a USNA research vessel in Chesapeake bay, Maryland.
The detailed configuration of the DCS composed of a base module and a rover module is presented at Figure 15. Specifically, the DCS includes Inertial Measurement Units (IMUs) and Relative Time Kinematic (RTK) GPS, from which an extended Kalman filter is executed to estimate the pose in real time. The rover module of the DCS has been enhanced with an Alvium 1800 C-240 Global Shutter RGB camera and a Jetson Nano single-board computer, connected via an MIPI CSI-2 cable. This setup allows image capture at a rate of 5Hz, synchronized with the RTK GPS via a WIFI connection. This upgraded rover configuration enables synchronization and time-stamping of the images within the RTK GPS data collection process. It allows for the effective collection of real image data along with accurate relative camera 6D poses.
@@
Figure 15: Data Collection System [8]
@@ 6.2 Model Testing with Real Data
The trained model is tested with the real world images captured by the DCS over multiple days. During the in-flight experiments conducted on the USNA research vessel, the model was subjected to real-world conditions where instances of variable lighting and occlusions were encountered. We deliberately selected challenging images for testing because they represent realistic conditions where such situations can occur. Figures 17 to 19 show the keypoints prediction and corresponding re-projected ships coordinate frame from estimated pose by the proposed TNN-MO model under the following three lighting conditions.
@@
Traditional feature extraction methods and visual marker-based approaches often struggle under challenging circumstances, as there is a lack of visual features on the ship. Furthermore, in this dynamic environment of real-world imaging, human operators and various items are present on the ship, which we can identify as occlusions. These occlusions, along with variable lighting, present significant challenges. However, as illustrated by Figures 17 to 19, the proposed model successfully identified the keypoints and the bounding boxes of multiple parts of the ship. Its robust performance under such conditions is a testament to its capability to provide accurate results, ensuring reliable 6D pose estimation even when faced with obstructions and fluctuating light levels.
Figure 17: Estimated keypoints and bounding boxes for overexposed images: objects with the confidence $o_{j}>0.9$ are highlighted
Figure 18: Estimated keypoints and bounding boxes for underexposed images: objects with the confidence $o_{j}>0.9$ are highlighted
Figure 19: Estimated keypoints and bounding boxes for normal images: objects with the confidence $o_{j}>0.9$ are highlighted
@@ 6.3 Validation with RTK GPS
Further, for quantitative validation, the estimated pose is compared with the actual pose determined by the DCS. The DCS integrates the relative attitude, derived from the base rover’s IMUs, and the relative position from the RTK-GPS with an extended Kalman filter. The IMUs, such as the VN-100, provide a magnetic heading accuracy of $2.0^{\circ}$, a gyro in-run bias stability of $5^{\circ}$/hr, and a pitch/roll accuracy of $0.5^{\circ}$ under normal conditions. The RTK-GPS, such as the Piksi Multi GNSS, provides a horizontal position accuracy of 0.75 m in SBAS mode, a velocity accuracy of 0.03 m/s RMS, and a time accuracy of 60 ns RMS. In RTK mode, it provides an accuracy of 0.010 m horizontally and 0.015 m vertically. These systems provide a centimeter-level accuracy in nominal conditions for the DCS. Here, the measurements of the RTK-GPS are considered as ground truth. The trajectories estimated by the proposed TNN-MO model and the RTK-GPS under the above three illumination conditions are illustrated in Figures 23 to 25 with the estimation errors and $3\sigma$-bounds calculated by (4) and (9), where it is shown that the position and the attitude trajectory estimated by the TNN-MO model are consistent with the IMU and the RTK-GPS.
@@
The mean absolute errors are also tabulated at Table 2. The MAE of position estimation ranges from $0.089$ to $0.177$ meters, which corresponds to $0.66\%$ to $0.97\%$ of the maximum range $L$. The MAE of rotation estimation varies from $1.1$ to $4.0$ degrees. While these errors are slightly larger than the validation error for the synthetic data as presented in Table 1, it is verified that the proposed TNN-MO successfully overcome the sim-to-real gap. The performance is consistent over the varying illumination conditions captured over multiple days. While the degree of attitude uncertainties measured by the standard deviation is in the reasonable range, the standard deviation for the position estimate is relatively small compared with the error, indicating over-confidence. This is partially because the uncertainties are measured indirectly in (4) by the variations of the estimate over multiple objects, not in the confidence of estimate for the individual object. However, in Figures 23 to 25, the uncertainties increase as the camera is further away from the ship, which is expected. These show that the presented uncertainty model behaves reasonably in the qualitative sense, but not necessarily accurate quantitatively. The exact modeling of uncertainties is considered as one of future directions. This provides promising results for pose estimate that can be utilized for autonomous launch and recovery in the ocean environments.
@@ 6.4 Multi-Objects Pose Estimation
Figure 20: Pose estimated with respect to the individual object class without fusion for underexposed images
For the second case of the underexposed ship, the trajectory estimated by each object is presented at Figure 20. Upon examining Figures 20(d) to 20(f), it is evident that when the camera is closer to the ship, three object categories of the super structure, the whole ship, and the stern are not within the field-of-view, and the corresponding estimate is degraded. Conversely, for Figures 20(a) to 20(c), object pose estimations are accurate and reliable, as the objects remain visible throughout the flight.
@@ 6.5 Attention Map
Figure 21: Encoder self-attention for a selected set of reference points: The attention map by using 300 ($H_{0}\times W_{0}$) features from the convoluted input image. Each pixel in the image corresponds to one of these features. We are visualizing how a feature linked to a selected pixel attends to other features. By analyzing the self-attention map, we can see that the encoder is capable of identifying key elements such as the sea, sky, and ship within the image based on the selected points.
Figure 22: Object keypoints are predicted by TNN-MO in the given real image (top), and the decoder cross-attention maps for the object queries corresponding to the predictions in the first row are shown (bottom): Here, we are visualizing how a encoder output attends to each queries related to the objects defined in Figure 9. However, the cross-attention related to $7$-th query, which is for the $\varnothing$ object class, is not presented in the figure.
To gain insights into how the synthetically trained TNN-MO model performs on real-world images, we analyzed the visualizations of the attention maps for the encoder and the decoder. The self-attention mechanism in the transformer network enables the encoder to understand the context of image elements. This is crucial in object detection, where the relationship between different parts of the image can provide important information. Therefore, we analyzed the transformer attention maps by passing a real-world image into TNN-MO, as shown in Figure 21. We observed that the TNN-MO encoder’s self-attention mechanism allows the model to understand the context of each pixel in relation to the others. For example, the self-attention map for the pixel point on the ship in the image, attention weights gets higher in the similar area of the ship on the map.
Additionally, to gain insights into what the model is focusing on, we visualized the decoder cross-attention map for different queries, as shown in Figure 22. This can help us understand which parts of the image the model finds most relevant for each query, providing valuable insights into the model’s decision-making process. We observed that the area where the cross-attention weight is higher for each query lies in the ship area of the image. This means that our TNN-MO model is able to understand the relevant information coming from the encoder to identify each part of the ship.

### v2

@@ 4.4 Model Inference with Bayesian Fusion
where $I_{0},I_{1}$ correspond to the modified Bessel function of the first kind, and $s\geq 0$ is a scalar that can be obtained numerically from $d$ [26]. The value of the above probability for varying $d$ and $\theta$ is illustrated at Fig. 11. For example, when $d=0.999$, the estimation error is less than $5.07^{\circ}$ with the probability of $0.95$.
Figure 11: Degree of confidence in the attitude estimate.
@@ 5.1 Model Training with Synthetic Data
The TNN-MO model was implemented using the PyTorch framework. The model was trained for 350 epochs with the batch size of 48, hyperparameter $\gamma=10$, leveraging the AdamW optimizer [27], using a 20GB Multi-Instance GPU (MIG) partition from NVIDIA A100-PCIE-40GB GPU. The learning rate was initially set to $10^{-4}$ for the first 200 epochs, after which it was reduced to $10^{-5}$. Gradient clipping was also employed, limiting the maximum gradient norm to 0.1. The resulting training curve is presented at Fig. 12.
Figure 12: Training loss over epochs.
@@ 5.2 Model Testing with Synthetic Data
Figure 13: TNN-MO model tested over synthetic data: for object classes with the confidence $o_{j}>0.9$, the key points and the base frame are illustrated in colors.
Figure 14: TNN-SO (top) vs TNN-MO (bottom) tested over long range synthetic data: for TNN-MO model, object classes with the confidence $o_{j}>0.9$, the key points are illustrated in colors.
@@ 6.1 Data Collection System (DCS)
The hardware configuration to capture a validation dataset is as follows. A data collection system (DCS) as shown in Fig. 16, that was developed for ship air wake measurements [15] is augmented with a camera. It is attached to an octocoper unmanned aerial vehicle that is manually controlled to fly around a USNA research vessel in Chesapeake bay, Maryland.
The detailed configuration of the DCS composed of a base module and a rover module is presented at Fig. 15. Specifically, the DCS includes Inertial Measurement Units (IMUs) and Relative Time Kinematic (RTK) GPS, from which an extended Kalman filter is executed to estimate the pose in real time. The rover module of the DCS has been enhanced with an Alvium 1800 C-240 Global Shutter RGB camera and a Jetson Nano single-board computer, connected via an MIPI CSI-2 cable. This setup allows image capture at a rate of 5Hz, synchronized with the RTK GPS via a WIFI connection. This upgraded rover configuration enables synchronization and time-stamping of the images within the RTK GPS data collection process. It allows for the effective collection of real image data along with accurate relative camera 6D poses.
@@
Figure 15: Data Collection System [13].
@@ 6.2 Model Testing with Real Data
The trained model is tested with the real world images captured by the DCS over multiple days. During the in-flight experiments conducted on the USNA research vessel, the model was subjected to real-world conditions where instances of variable lighting and occlusions were encountered. We deliberately selected challenging images for testing because they represent realistic conditions where such situations can occur. Figs. 17 to 19 show the keypoints prediction and corresponding re-projected ships coordinate frame from estimated pose by the proposed TNN-MO model under the following three lighting conditions.
@@
Traditional feature extraction methods and visual marker-based approaches often struggle under challenging circumstances, as there is a lack of visual features on the ship.
Figure 17: Estimated keypoints and bounding boxes for overexposed images: objects with the confidence $o_{j}>0.9$ are highlighted.
Figure 18: Estimated keypoints and bounding boxes for underexposed images: objects with the confidence $o_{j}>0.9$ are highlighted.
Figure 19: Estimated keypoints and bounding boxes for normal images: objects with the confidence $o_{j}>0.9$ are highlighted.
Furthermore, in this dynamic environment of real-world imaging, human operators and various items are present on the ship, which we can identify as occlusions. These occlusions, along with variable lighting, present significant challenges. However, as illustrated by Figs. 17 to 19, the proposed model successfully identified the keypoints and the bounding boxes of multiple parts of the ship. Its robust performance under such conditions is a testament to its capability to provide accurate results, ensuring reliable 6D pose estimation even when faced with obstructions and fluctuating light levels.
@@ 6.3 Validation with RTK GPS
Further, for quantitative validation, the estimated pose is compared with the actual pose determined by the DCS. The DCS integrates the relative attitude, derived from the base rover’s IMUs, and the relative position from the RTK-GPS with an extended Kalman filter. The IMUs, such as the VN-100, provide a magnetic heading accuracy of $2.0^{\circ}$, a gyro in-run bias stability of $5^{\circ}$/hr, and a pitch/roll accuracy of $0.5^{\circ}$ under normal conditions. The RTK-GPS, such as the Piksi Multi GNSS, provides a horizontal position accuracy of 0.75 m in SBAS mode, a velocity accuracy of 0.03 m/s RMS, and a time accuracy of 60 ns RMS. In RTK mode, it provides an accuracy of 0.010 m horizontally and 0.015 m vertically. These systems provide a centimeter-level accuracy in nominal conditions for the DCS. Here, the measurements of the RTK-GPS are considered as ground truth. The trajectories estimated by the proposed TNN-MO model and the RTK-GPS under the above three illumination conditions are illustrated in Figs. 20 to 22 with the estimation errors and $3\sigma$-bounds calculated by (4) and (9), where it is shown that the position and the attitude trajectory estimated by the TNN-MO model are consistent with the IMU and the RTK-GPS.
(a) 3d view of the ship
(b) Top and side view of the ship
(c) Position, $x$
(d) Position Error, $e_{x}$
(e) Attitude, $ypr$
(f) Attitude Error, $\eta$
Figure 20: Overexposed case: the estimated pose (red) is compared against the RTK-GPS measurements (blue).
(a) 3d view of the ship
(b) Top and side view of the ship
(c) Position, $x$
(d) Position Error, $e_{x}$
(e) Attitude, $ypr$
(f) Attitude Error, $\eta$
Figure 21: Underexposed case: the estimated pose (red) is compared against the RTK-GPS measurements (blue).
(a) 3d view of the ship
(b) Top and side view of the ship
(c) Position, $x$
(d) Position Error, $e_{x}$
(e) Attitude, $ypr$
(f) Attitude Error, $\eta$
Figure 22: Nominal case: the estimated pose (red) is compared against the RTK-GPS measurements (blue).
@@
The mean absolute errors are also tabulated at Table 2. The MAE of position estimation ranges from $0.089$ to $0.177$ meters, which corresponds to $0.66\%$ to $0.97\%$ of the maximum range $L$. The MAE of rotation estimation varies from $1.1$ to $4.0$ degrees. While these errors are slightly larger than the validation error for the synthetic data as presented in Table 1, it is verified that the proposed TNN-MO successfully overcome the sim-to-real gap. The performance is consistent over the varying illumination conditions captured over multiple days. While the degree of attitude uncertainties measured by the standard deviation is in the reasonable range, the standard deviation for the position estimate is relatively small compared with the error, indicating over-confidence. This is partially because the uncertainties are measured indirectly in (4) by the variations of the estimate over multiple objects, not in the confidence of estimate for the individual object. However, in Figs. 23 to 25, the uncertainties increase as the camera is further away from the ship, which is expected. These show that the presented uncertainty model behaves reasonably in the qualitative sense, but not necessarily accurate quantitatively. The exact modeling of uncertainties is considered as one of future directions. This provides promising results for pose estimate that can be utilized for autonomous launch and recovery in the ocean environments.
@@ 6.4 Multi-Objects Pose Estimation
Figure 23: Pose estimated with respect to the individual object class without fusion for underexposed images.
For the second case of the underexposed ship, the trajectory estimated by each object is presented at Fig. 23. Upon examining Figs. 23(d) to 23(f), it is evident that when the camera is closer to the ship, three object categories of the super structure, the whole ship, and the stern are not within the field-of-view, and the corresponding estimate is degraded. Conversely, for Figs. 23(a) to 23(c), object pose estimations are accurate and reliable, as the objects remain visible throughout the flight.
@@ 6.5 Attention Map
Figure 24: Encoder self-attention for a selected set of reference points: The attention map by using 300 ($H_{0}\times W_{0}$) features from the convoluted input image.
To gain insights into how the synthetically trained TNN-MO model performs on real-world images, we analyzed the visualizations of the attention maps for the encoder and the decoder. The self-attention mechanism in the transformer network enables the encoder to understand the context of image elements. This is crucial in object detection, where the relationship between different parts of the image can provide important information. Therefore, we analyzed the transformer self-attention maps by passing a real-world image into TNN-MO. In Fig. 24, the encoder’s self-attention for a selected set of reference points is visualized in the attention map using 300 ($H_{0}\times W_{0}$) features from the convoluted input image. Each pixel in the image corresponds to one of these features. We are visualizing how a feature linked to a selected pixel attends to other features. By analyzing the self-attention map, we can see that the encoder is capable of identifying key elements such as the sea, sky, and ship within the image based on the selected points. We observed that the TNN-MO encoder’s self-attention mechanism allows the model to understand the context of each pixel in relation to the others. For example, the self-attention map for the pixel point on the ship in the image, attention weights gets higher in the similar area of the ship on the map.
Figure 25: Object keypoints are predicted by TNN-MO in the given real image (top), and the decoder cross-attention maps for the object queries corresponding to the predictions in the first row are shown (bottom).
Additionally, to gain insights into what the model is focusing on, we visualized the decoder cross-attention map for different queries. In Fig. 25, we are visualizing how a encoder output attends to each queries related to the objects defined in Fig. 9, excluding the cross-attention related to $7$-th query, which is for the $\varnothing$ object class. This can help us understand which parts of the image the model finds most relevant for each query, providing valuable insights into the model’s decision-making process. We observed that the area where the cross-attention weight is higher for each query lies in the ship area of the image. This means that our TNN-MO model is able to understand the relevant information coming from the encoder to identify each part of the ship.